How many grams of CO2 are produced when 17.0 gal of C8H18 are combusted
Answer #1
2C8H18 + 25O2 = 16CO2 + 18 H2O

17 gal =17*3.7854l =64.3518 l

Moles of octane = 64.3518/22 =2.925 mol

Here 2 moles produce 16 mole of gas
So 2.925 moles produce = 16*2.925/2 =23.4 mol
Mass =23.4*44 =1029.6288 gram
answered by: Patty Ann